OpenStudy (anonymous):

Determine the sum of the measures of the interior angles of the specified polygon. A dodecagon A. 1440° B. 1620° C. 2160° D. 1800°

OpenStudy (anonymous):

the equation is (10 - 2) * 180 = 8 * 180 = 1440

OpenStudy (anonymous):

the general equation being (n - 2) * 180

OpenStudy (experimentx):

1800

OpenStudy (anonymous):

http://answers.yahoo.com/question/index?qid=20100611120848AA0JvbQ the majority say 1440 on yahoo answers

OpenStudy (experimentx):

dodecagon has 12 sides